OpenStudy (anonymous):

Determine the Taylor series to represent cos(pi/3+h). Hence evaluate cos60.001 to 8 decimal places reply soon

OpenStudy (dumbcow):

let h =0 f(h) = 1/2 f'(h) = -sqrt3/2 f''(h) = -1/2 f'''(h) = sqrt3/2 then it repeats \[\cos(\pi/3 + h) = \frac{1}{2}-\frac{\sqrt{3}}{2}h-\frac{1}{4}h^{2}+\frac{\sqrt{3}}{12}h^{3}+\frac{1}{48}h^{4} ...\]

OpenStudy (dumbcow):

you will prob need to add 3 or 4 more terms to get the accuracy you need